# Find the Fruit: Zero-Shot Sim2Real RL for Occlusion-Aware Plant Manipulation

### v2

@@ I Introduction
Autonomous harvesting robots must do more than grasping a fruit — they must first find it. In most real world scenarios, fruits are generally hidden deep within dense foliage, partially or completely occluded by leaves, stems, or branches. These occlusions slow down harvesting, increase missed detection rates, and often require manual repositioning before automated grasping can even begin. While recent advances have led to robust perception for visible fruits and reliable manipulation for grasping, the problem of occlusion-resolution for harvesting remains underexplored. Most existing harvesting systems either assume fruits are already visible [1, 2, 3], or adjust the robot itself to get a clearer view of the fruit [4, 5, 6, 7]. However, such strategies fail to scale in real scenarios, where visibility could vary with growth stage, fruit distribution, and foliage density. For robust harvesting, a robot needs to actively manipulate the plant to reveal fruits before grasping — we refer it as occlusion-aware fruit discovery. Addressing this problem is essential for enabling downstream systems to function robustly in real-world conditions.
@@
Perception for harvesting is challenging due to significant occlusions: leaves, stems, and branches often hide fruits partially or completely, breaking the assumptions of standard detection pipelines [8, 9, 10]. Planning is complicated by cluttered and variable plant geometry [11, 12]. Interaction during manipulation is particularly difficult due to deformable structure of plants. Flexible stems bend, leaves fold, and branches recoil unpredictably under contact. Designing controllers in such settings is difficult since stiffness, damping, and mass distribution vary across species and even between individuals [13, 14]. These uncertainties make open-loop strategies and model-based control brittle, especially when occlusions change dynamically during interaction.
@@
Existing research on deformable object manipulation does not fully transfer to agricultural settings. Prior work has largely focused on ropes and chains [15, 16, 17], fabrics [18, 19, 20, 21], or biological tissues [22], often under controlled conditions with strong modeling assumptions. Agricultural studies have explored leaf removal [9, 23], navigating through the branches to reach parts of the plant [24], sequential exploration to predict newly revealed spaces [25], or branch repositioning prior to harvesting [26, 27, 28]. Other efforts target collision-free grasp planning [29, 30]. While valuable, many of these methods depend on explicit detection of plant parts or detailed 3D reconstructions [31, 32, 33], which are computationally expensive and difficult to scale to highly variable and deformable foliage.
@@
We believe that a closed-loop learning approach offers a promising alternative. Since occlusion patterns can change during manipulation, traditional motion planning-based solutions tend to struggle [34, 35, 36]. Reinforcement Learning (RL) can adaptively couple perception and action. By leveraging sensory feedback, RL agents can react to plant deformation without requiring brittle intermediate models. However, sim2real transfer is a significant challenge  [37, 38, 39], which is especially pronounced for deformable plants because their mechanical properties are difficult to model faithfully.
@@
Our Approach: We propose an end-to-end RL framework for occlusion-aware manipulation of deformable plants with zero-shot sim-to-real transfer (see Fig. LABEL:fig:abstract_figure). The policy is trained entirely in simulation using a generic plant model built in NVIDIA Isaac Lab [40], with deformable dynamics and randomized occlusion patterns. In our simulation setup, the plant is represented using a finite element method (FEM) model rather than simplified kinematic chains or reduced-order approximations. This formulation directly captures bending, twisting, and nonlinear deformation of stems and leaves under contact, without relying on lumped stiffness parameters or geometric heuristics. By avoiding lower-order assumptions, the model preserves the continuous deformation behavior of foliage, providing the RL policy with interaction dynamics that are structurally faithful to real plants [13, 41]. This high-fidelity representation serves as a key inductive bias, enabling the learned policy to transfer more reliably to physical systems. This model is used to train the policy entire in simulation using an on-policy algorithm (PPO) [42].
@@ II-E Deployment Framework
The deployment framework, as shown in Fig. 2, integrates the trained RL controller with a low-level servo position controller. The end-effector pose and joint angles are published to ROS 2, where they are subscribed to by the control PC. Simultaneously, an RGB-D image is captured from the image sensor. These observations are fed into the policy, which subsequently generates an action. The resulting action, converted to the target joint angles, is then published and subscribed to by the robot. At this stage, only kinematic control is involved. Consequently, the servo position controller attempts to reach the target position in a controlled manner without exerting excessive force on the plant. This process operates in a continuous loop, maintaining safe and effective control.
@@ IV Results
This section presents the outcomes of deploying our learned policy on our real setup, focusing on the observed behaviors, quantitative success rates, and identified failure modes and limitations. We first analyze policy behavior and generalization across different plant morphologies, followed by an extended evaluation to sequential multi-fruit scenarios, then ablation studies on the role of privileged information and simulation fidelity. Finally, we identify and discuss the failure modes. Together, these results provide a picture of how the proposed framework performs under real-world conditions.
@@ IV-B Multiple Fruit Generalization
Our policy is trained primarily in a single-fruit setup. Nevertheless, it could generalizes to multiple targets without architectural changes by following a sequential exposure protocol: once one fruit is uncovered, it is removed from the scene, and the same policy continues searching for the remaining fruit(s). This removal reflects a practical harvesting scenario in which one arm exposes the fruit while another arm can perform the picking task. Importantly, the control loop runs continuously, we do not reset the robot or environment after the first fruit is found. In scenarios where the first fruit is not to be harvested, e.g., unripe, though not implemented in our experiments, we could apply an ignore mask in the observation stream to prevent re-targeting while proceeding to the next fruit. The testing procedures are as follows:
@@ VI Limitations
(a) Simplified simulation: Training relied on a single-stem, single-fruit abstraction. While sufficient for stemmed plants, it does not capture more complex morphologies such as multi-branch canopies or fruit clusters. We also assumed that fruit state remains unchanged during plant manipulation and that perception comes from a fixed overhead camera rather than a robot-mounted sensor that can move during the operation. Extending the simulator to relax these assumptions and include richer structures is a natural next step.
@@
VII Acknowledgment
This work is jointly funded from NSF-USDA COALESCE grant #2021-67021-34418 and AIIRA grant #2021-67021-35329

### v3

@@ I Introduction
Autonomous harvesting robots must do more than grasping a fruit – they must first find it. In most real world scenarios, fruits are generally hidden deep within dense foliage, partially or completely occluded by leaves, stems, or branches. These occlusions slow down harvesting, increase missed detection rates, and often require manual repositioning before automated grasping can even begin. While recent advances have led to robust perception for visible fruits and reliable manipulation for grasping, the problem of occlusion-resolution for harvesting remains underexplored. Most existing harvesting systems either assume fruits are already visible [36, 1, 39], or adjust the robot itself to get a clearer view of the fruit [19, 13, 2, 41]. However, such strategies fail to scale in real scenarios, where visibility could vary with growth stage, fruit distribution, and foliage density. For robust harvesting, a robot needs to actively manipulate the plant to reveal fruits before grasping — we refer it as occlusion-aware fruit discovery. Addressing this problem is essential for enabling downstream systems to function robustly in real-world conditions.
@@
Perception for harvesting is challenging due to significant occlusions: leaves, stems, and branches often hide fruits partially or completely, breaking the assumptions of standard detection pipelines [5, 44, 14]. Planning is complicated by cluttered and variable plant geometry [26, 34]. Interaction during manipulation is difficult due to deformable structure of plants. Flexible stems bend, leaves fold, and branches recoil unpredictably under contact. Designing controllers in such settings is difficult since stiffness, damping, and mass distribution vary across species and even between individuals [6, 30]. These uncertainties make open-loop strategies and model-based control brittle, especially when occlusions change dynamically during interaction.
@@
Existing research on deformable object manipulation does not fully transfer to agricultural settings. Prior work has largely focused on ropes and chains [4, 25, 22], fabrics [7, 8, 20, 3], or biological tissues [31], often under controlled conditions with strong modeling assumptions. Agricultural studies have explored leaf removal [44, 43], navigating through the branches to reach parts of the plant [10], sequential exploration to predict newly revealed spaces [45], or branch repositioning prior to harvesting [16, 11, 29]. Other efforts target collision-free grasp planning [32, 37]. While valuable, many of these methods rely on explicit plant-part detection or detailed 3D reconstruction [27, 12, 23], which are computationally costly and hard to scale to highly deformable, variable foliage.
@@
We believe that a closed-loop learning approach offers a promising alternative. As occlusions evolve during manipulation, classical motion planning methods often struggle [17, 9, 21]. Reinforcement learning (RL) adaptively couples perception and action, enabling agents to respond to plant deformation without brittle intermediate models. However, sim-to-real transfer remains challenging [40, 38, 46], particularly for deformable objects with difficult-to-model mechanics [42].
@@
Our Approach: We propose an end-to-end RL framework for occlusion-aware manipulation of deformable plants with zero-shot sim-to-real transfer (see Fig. LABEL:fig:abstract_figure). The policy is trained entirely in simulation using a generic plant model built in NVIDIA Isaac Lab [24], with deformable dynamics and randomized occlusion patterns. In our simulation setup, the plant is represented using a finite element method (FEM) model rather than simplified kinematic chains or reduced-order approximations. This formulation directly captures bending, twisting, and nonlinear deformation of stems and leaves under contact, without relying on lumped stiffness parameters or geometric heuristics. By avoiding lower-order assumptions, the model preserves the continuous deformation behavior of foliage, providing the RL policy with interaction dynamics that are structurally faithful to real plants [6, 28]. This high-fidelity representation serves as a key inductive bias, enabling the learned policy to transfer more reliably to physical systems [18]. This model trains the policy entirely in simulation using PPO [33].
@@ II-E Deployment Framework
The deployment framework (Fig. 2) integrates the trained RL controller with a low-level servo position controller. The end-effector pose and joint angles are published to ROS 2 and subscribed to by the control PC. Simultaneously, an RGB-D image is captured from the sensor. These observations are fed into the policy, which generates an action. The action, converted to target joint angles, is then published and received by the robot. At this stage, only kinematic control is used. The servo controller tracks the target position in a controlled manner without exerting excessive force on the plant. This process runs in a continuous loop, ensuring safe and effective control.
@@ IV Results
This section presents results from deploying our learned policy on the real setup, focusing on observed behaviors, success rates, and failure modes. We first analyze generalization across plant morphologies, followed by evaluation in sequential multi-fruit scenarios and ablations on privileged information and simulation fidelity. Finally, we discuss key failure cases and limitations.
@@ IV-B Multiple Fruit Generalization
Our policy is trained primarily in a single-fruit setup. Nevertheless, it could generalizes to multiple targets without architectural changes by following a sequential exposure protocol: once one fruit is uncovered, it is removed from the scene, and the same policy continues searching for the remaining fruit(s). This removal reflects a practical harvesting scenario in which one arm exposes the fruit while another arm can perform the picking task. Importantly, the control loop runs continuously, we do not reset the robot or environment after the first fruit is found. In scenarios where the first fruit is not to be harvested, e.g., unripe, though not implemented in our experiments, we could apply an ignore mask in the observation stream to prevent re-targeting while proceeding to the next fruit. The testing procedures are:
@@ VI Limitations
(a) Simplified simulation: Training relied on a single-stem, single-fruit abstraction. While adequate for simple plants, it does not capture complex morphologies like multi-branch canopies or fruit clusters. We also assumed static fruit state and a fixed overhead camera. Extending the simulator to relax these assumptions is a natural next step.
@@
Acknowledgment
This work is jointly funded by the NSF–USDA COALESCE program (NSF award #1954556; USDA award #2021-6702134418)